OpenStudy (anonymous):

PLEASE HELP AWARDING METAL!! Verify each trigonometric equation by substituting identities to match the right hand side of the equation to the left hand side of the equation. 1 + sec2x sin2x = sec2x

OpenStudy (michele_laino):

left side: we can replace sec x with 1/cos x so we can write: \[1 + {\left( {\sec x} \right)^2}{\left( {\sin x} \right)^2} = 1 + \frac{{{{\left( {\sin x} \right)}^2}}}{{{{\left( {\cos x} \right)}^2}}}\]

OpenStudy (anonymous):

How?

OpenStudy (michele_laino):

here are more steps: \[1 + {\left( {\sec x} \right)^2} \times {\left( {\sin x} \right)^2} = 1 + \frac{1}{{{{\left( {\cos x} \right)}^2}}} \times {\left( {\sin x} \right)^2} = 1 + \frac{{{{\left( {\sin x} \right)}^2}}}{{{{\left( {\cos x} \right)}^2}}}\]

OpenStudy (anonymous):

what is the first step?

OpenStudy (michele_laino):

now we can write this: \[1 + \frac{{{{\left( {\sin x} \right)}^2}}}{{{{\left( {\cos x} \right)}^2}}} = \frac{{{{\left( {\cos x} \right)}^2} + {{\left( {\sin x} \right)}^2}}}{{{{\left( {\cos x} \right)}^2}}} = ...?\]

OpenStudy (michele_laino):

please simplify

OpenStudy (anonymous):

I don't have much time!

OpenStudy (michele_laino):

you have to use the fundamental identity of trigonometry, namely: \[{\left( {\cos x} \right)^2} + {\left( {\sin x} \right)^2} = 1\]

OpenStudy (michele_laino):

so we can write: \[1 + \frac{{{{\left( {\sin x} \right)}^2}}}{{{{\left( {\cos x} \right)}^2}}} = \frac{{{{\left( {\cos x} \right)}^2} + {{\left( {\sin x} \right)}^2}}}{{{{\left( {\cos x} \right)}^2}}} = \frac{1}{{{{\left( {\cos x} \right)}^2}}}\]

OpenStudy (anonymous):

ok this is the first step...

OpenStudy (michele_laino):

right side: we can rewrite the right side, as below: \[{\left( {\sec x} \right)^2} = \frac{1}{{{{\left( {\cos x} \right)}^2}}}\]

OpenStudy (michele_laino):

so we have shown that left side is equal to right side, and our work is completed

OpenStudy (anonymous):

same problem, different question. - tan2x + sec2x = 1

OpenStudy (michele_laino):

here we have to replace tan x with its definition, namely tan x = sin x/ cos x, and we have to replace sec x with its definition, namely: sec x= 1/ cos x, so we get: \[ - {\left( {\tan x} \right)^2} + {\left( {\sec x} \right)^2} = - \frac{{{{\left( {\sin x} \right)}^2}}}{{{{\left( {\cos x} \right)}^2}}} + \frac{1}{{{{\left( {\cos x} \right)}^2}}} = ...?\] please continue

OpenStudy (anonymous):

please give me one minute.

OpenStudy (michele_laino):

ok!

OpenStudy (anonymous):

Thank you so much for your Patience :)

OpenStudy (michele_laino):

thanks! :)

OpenStudy (anonymous):

could you go over it and explain after.

OpenStudy (michele_laino):

yes I can, nevertheless you also have to contribute in getting your answers

OpenStudy (michele_laino):

next step is: \[ - \frac{{{{\left( {\sin x} \right)}^2}}}{{{{\left( {\cos x} \right)}^2}}} + \frac{1}{{{{\left( {\cos x} \right)}^2}}} = \frac{{1 - {{\left( {\sin x} \right)}^2}}}{{{{\left( {\cos x} \right)}^2}}} = ...?\]

OpenStudy (anonymous):

Ok, sorry time is just working against me here.

OpenStudy (michele_laino):

please keep in mind that I have to respect the Code of Conduct

OpenStudy (michele_laino):

now we have: \[ - \frac{{{{\left( {\sin x} \right)}^2}}}{{{{\left( {\cos x} \right)}^2}}} + \frac{1}{{{{\left( {\cos x} \right)}^2}}} = \frac{{1 - {{\left( {\sin x} \right)}^2}}}{{{{\left( {\cos x} \right)}^2}}} = \frac{{{{\left( {\cos x} \right)}^2}}}{{{{\left( {\cos x} \right)}^2}}} = ...?\]

OpenStudy (anonymous):

Do I simplify?

OpenStudy (michele_laino):

yes!

OpenStudy (michele_laino):

hint: 

OpenStudy (anonymous):

−(sinx)2/(cosx)2+1/(cosx)2=1−(sinx)2/(cosx)2=

OpenStudy (michele_laino):

yes! and the next step is: \[\frac{{1 - {{\left( {\sin x} \right)}^2}}}{{{{\left( {\cos x} \right)}^2}}} = \frac{{{{\left( {\cos x} \right)}^2}}}{{{{\left( {\cos x} \right)}^2}}} = ...?\]

OpenStudy (michele_laino):

so, what do you get?

OpenStudy (anonymous):

now we have: −(sinx)2/(cosx)2+1/(cosx)2=1−(sinx)2/(cosx)2= 1-(sinx)^2/ (cosx)^2

OpenStudy (anonymous):

Right?

OpenStudy (michele_laino):

yes! it is right! nevertheless you have to simplify this expression: \[\frac{{{{\left( {\cos x} \right)}^2}}}{{{{\left( {\cos x} \right)}^2}}} = ...?\]

OpenStudy (michele_laino):

is the result 1?

OpenStudy (anonymous):

yes

OpenStudy (michele_laino):

ok! so we have shown that the left side is equal to 1, which is the right side, then our work is completed

OpenStudy (michele_laino):

please wait a moment I have to answer to my phone

OpenStudy (anonymous):

Ok.

OpenStudy (anonymous):

1*1/ (sinx)^2/ (cosx)^2

OpenStudy (michele_laino):

here I am

OpenStudy (anonymous):

Right?

OpenStudy (michele_laino):

can you write your expression with the editor, since I don't understand it

OpenStudy (michele_laino):

or please make a drawing of your expression

OpenStudy (anonymous):



OpenStudy (michele_laino):

is your expression like this? \[\Large \frac{{1 \times 1}}{{{{\left( {\sin x} \right)}^2}{{\left( {\cos x} \right)}^2}}}\]

OpenStudy (anonymous):

Yes, is it right?

OpenStudy (michele_laino):

are you referring to the second exercise?

OpenStudy (anonymous):

nevermind...

OpenStudy (michele_laino):

the first two identities have been checked have you another exercise to solve?

OpenStudy (anonymous):

yes! one last one. Give me one minute.

OpenStudy (michele_laino):

ok!

OpenStudy (michele_laino):

@clairvoyant1

OpenStudy (anonymous):

hey sorry, let us continue.

OpenStudy (michele_laino):

ok!

OpenStudy (michele_laino):

what is your question?

OpenStudy (michele_laino):

@clairvoyant1

OpenStudy (michele_laino):

Please, if you need help with your mathematics, then tag me @clairvoyant1

OpenStudy (anonymous):

\[\frac{ sinx }{ 1-cosx }+\frac{ sinx }{ 1+cosx }=2 \csc x\]

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

@welshfella